OpenStudy (anonymous):

sqrt(2x+1)-sqrt(x)=1. Solve for x.

OpenStudy (anonymous):

x=0

OpenStudy (anonymous):

x = 0 and x = 4

OpenStudy (anonymous):

first square both sides u will get 2x + 1 + x + 2 ( sqrt x(2x+1)) = 1 so it will be 3x = -2 ( sqrt x(2x+1)) again square on both sides 9 x^2 = 4x(2x + 1) u willl get 9x = 8x + 4 x = 4

OpenStudy (anonymous):

since will cancel out x between the steps u can consider x =0 also

OpenStudy (anonymous):

the answer is 0 and 4. how can i solve?

OpenStudy (anonymous):

hey i gave u the solution right?

OpenStudy (nikvist):

\[\sqrt{2x+1}-\sqrt{x}=1\quad,\quad x\geq 0\]\[\sqrt{2x+1}=\sqrt{x}+1\]\[2x+1=x+1+2\sqrt{x}\]\[x=2\sqrt{x}\]\[x(x-4)=0\quad,x=0\vee x=4\]